OpenStudy (anonymous):

use the angle sum identity to find the exact value of each: cos 105°

OpenStudy (anonymous):

You need to know the equation for cos(a+b)=? and then know what cos(30) cos(60) cos(45) and probably the sin()s as well.

OpenStudy (anonymous):

cos(a+b) = cos(a)cos(b)-sin(a)sin(b) so if you take 105=60 + 45 then you have cos(60+45) = cos(60)cos(45)-sin(60)sin(45)